# Revealing Fine-Grained Values and Opinions in Large Language Models

### v1

@@ Abstract
Uncovering latent values and opinions in large language models (LLMs) can help identify biases and mitigate potential harm. Recently, this has been approached by presenting LLMs with survey questions and quantifying their stances towards morally and politically charged statements. However, the stances generated by LLMs can vary greatly depending on how they are prompted, and there are many ways to argue for or against a given position. In this work, we propose to address this by analysing a large and robust dataset of 156k LLM responses to the 62 propositions of the Political Compass Test (PCT) generated by 6 LLMs using 420 prompt variations. We perform coarse-grained analysis of their generated stances and fine-grained analysis of the plain text justifications for those stances. For fine-grained analysis, we propose to identify tropes in the responses: semantically similar phrases that are recurrent and consistent across different prompts, revealing patterns in the text that a given LLM is prone to produce. We find that demographic features added to prompts significantly affect outcomes on the PCT, reflecting bias, as well as disparities between the results of tests when eliciting closed-form vs. open domain responses. Additionally, patterns in the plain text rationales via tropes show that similar justifications are repeatedly generated across models and prompts even with disparate stances.
Revealing Fine-Grained Values and Opinions in Large Language Models
Dustin Wrightthanks:   denotes equal contribution      Arnav Arora11footnotemark: 1      Nadav Borenstein      Srishti Yadav Serge Belongie      Isabelle Augenstein University of Copenhagen & Pioneer Centre for AI {dw, aar, nb, srya, s.belongie, augenstein}@di.ku.dk
@@ 1 Introduction
First, recent work has shown that the responses of LLMs to survey questions depend highly on the phrasing of the question and the format of the answer (Wang et al., 2024; Röttger et al., 2024; Motoki et al., 2024), calling for a more robust evaluation setup for surfacing values embedded in language models. Second, when provided with different personas based on demographic characteristics, LLMs can reflect the social and political biases of the respective demographics Argyle et al. (2023), highlighting the need for disentangling the opinions embedded into LLMs and their variation when prompted with demographics. Such efforts also aid in aligning language models for different populations and cultures and prevent jailbreaking of LLMs Wei et al. (2023). Lastly, these evaluations focus primarily on quantifying stances towards the survey questions, ignoring justifications and explanations for those decisions. Revealing biases in such data could further demonstrate latent values and opinions as expressed in plain text.
To address these shortcomings, we conduct a large-scale study eliciting 156,240 responses to the 62 propositions of the Political Compass Test (PCT) across 6 LLMs and 420 prompt variations. These prompt variations span different demographic personas: Age, Gender, Nationality, Political Orientation, and Class, as well as instruction prompts, in order to provide a robust set of data for analysis while disentangling demographic features. We propose to perform both coarse-grained analysis at the level of stances, allowing us to quantify political bias based on the PCT, as well as fine-grained analysis of the plain text open-ended justifications and explanations for those stances, allowing us to reveal latent values and opinions in the generated text. For fine-grained analysis, we propose to identify tropes in the responses: semantically similar phrases that are recurrent and consistent across different prompts, revealing patterns in the justifications that LLMs are prone to produce in different settings. An example from our dataset is given in Fig. 1, where Llama 3 AI@Meta (2024) when prompted with two different demographic personas (being far left vs. far right) demonstrates both a tendency towards particular stances as well as particular lines of reasoning for those stances.
@@
  • We propose a new method for analysing bias in generated text through tropes, revealing the arguments which LLMs are likely to generate across prompts in different settings;
@@ 2 Related Work
Surfacing political biases embedded in NLP tools has been approached previously in the context of word embeddings Gordon et al. (2020) and masked language modeling (Schramowski et al., 2022; Stańczak et al., 2023). As these biases originate in the data the models are trained on Borenstein et al. (2024), there is also research on directly tying them back to training data (Feng et al., 2023). Recently, with more performant and coherent LLMs, extracting inherent biases have garnered more attention. This has been explored by eliciting responses from psychological surveys Miotto et al. (2022), opinion surveys Arora et al. (2023); Durmus et al. (2023), and personality tests (Rutinowski et al., 2023), finding that certain LLMs have a left-libertarian bias (Hartmann et al., 2023). There has also been research on examining the framing of the generations produced by LLMs (Bang et al., 2024), and on measuring biases across different social bias categories (Manerba et al., 2024). Recently, persona based evaluation of political biases has also been attempted to see if models can simulate the responses of populations (Liu et al., 2024; Hu and Collier, 2024; Jiang et al., 2022), including in languages other than English Thapa et al. (2023). However, these remain underexplored areas.
Furthermore, there is research demonstrating the brittleness, values, and opinions of LLMs (Ceron et al., 2024). Our work is most similar to Röttger et al. (2024), who try to answer the question of how to meaningfully evaluate values and opinions in LLMs by way of the PCT. They demonstrate that eliciting open-ended responses from LLMs can lead to vastly different results than closed-form categorical selections, motivating the use of open-ended responses. We take this a step further, demonstrating the systematic variance of LLM responses to the PCT across different types of demographic prompts and further highlighting biased values and opinions in the open-ended responses.
@@ 3.1 Robust Dataset Generation
We generate a large dataset of responses to the PCT by varying two aspects of the prompt for each proposition: one aspect controlling the demographic category, and one aspect controlling the generation instruction. We improve over previous work, which has looked at robustness by varying the instructions or the order of appearance of options in the prompt Röttger et al. (2024); Hartmann et al. (2023); Motoki et al. (2024), by also varying across demographic features and analyzing the plain text justifications for the provided stances. As such, we adopt a general prompt template of the following form which we subsequently fill for the response to be generated: Prompt Template: Please follow all instructions from the perspective of a person who is {demographic}. Please write all responses in English. {instruction} where the demographic is selected from Table 1 and the instruction is selected from either Table 3 or Table 4 (see Appendix A.1). We use 21 different demographic options, 20 instruction variations, and there are 62 PCT propositions, resulting in 26,040 responses per model that we study.
We cover both breadth and depth of demographic categories, covering 5 types of demographics and 3-5 values for each. For the instruction, prior work has shown that responses of an LLM change when constraints are put on the answer generation format (Röttger et al., 2024). Therefore, we use two settings: 1) open-ended generation, where no constraint is put on the model in terms of choosing a particular option; and 2) closed form generation, where the model is explicitly prompted to choose one of the listed options. In the closed form generation setting, the model is prompted to choose a stance towards the position based on the listed responses in the PCT survey: Strongly Agree, Agree, Disagree, Strongly Disagree. Additionally, we prompt the model to output an explanation for the selection. In the open-ended generation setting, the model is prompted with an open ended-prompt with no additional constraints. To further conduct coarse-grained analysis of the responses from the open-ended generation setting, including alignment between the open and closed settings, we categorise the open-ended responses of the LLMs into the selection options from the closed setting post-hoc using a Mistral-Instruct-v0.2 model (79% accuracy on a held out test set manually annotated by 3 annotators, see Appendix A.6 for more details).
@@ 3.2 Tropes Extraction
Measuring categorical stances generated towards the PCT propositions provides coarse grained information about values and opinions by allowing one to quantify the political alignment of a given model/prompt combination. However, this coarse-grained information disregards the plain text justifications and explanations a model is likely to generate with respect to the propositions, which may reveal latent values and opinions not measured by the stances. In practice, users interact with LLMs in a plain-text, open-ended fashion, which this fine-grained information reflects. As such, we propose to complement the use of categorical stances with analysis of tropes present in the plain text responses. [^4] Tropes have been well studied in multiple domains, including literature Miller (1991) and television Gala et al. (2020). We adopt the following definition of a trope: A theme or motif which has the following properties: It is recurrent, i.e., appears frequently in the responses; and it is consistent, i.e., the statements which represent the trope can be grounded in a single abstract concept. Additionally, we focus specifically on tropes which convey a justification or explanation for the stances generated toward the political propositions. An example of such tropes is given in Fig. 1. To find tropes, we propose to: 1) generate many responses for each proposition under different conditions (§3.1); 2) cluster individual sentences based on their semantic similarity, and (§3.2.1); 3) distill the semantic clusters to single high-level concepts, filtering those clusters which do not contain justifications or explanations relevant to the stance (§3.2.2).
@@ 3.2.2 Distilling Tropes
[…]
@@ 4 Analysis
Our analysis centers around four research questions addressing the three shortcomings of previous work described in §1: RQ1: How do demographic-based prompts impact LLM survey responses? RQ2: Do categorical surveys administered to LLMs elicit robust results over diverse prompts? RQ3: What tropes do LLMs produce in response to the PCT? RQ4: Do variations in categorical survey responses reflect variations in the tropes? To answer these questions, we generate a large dataset of responses to the PCT as described in §3.1, eliciting 26,040 responses for 6 different language models (156,240 responses total). [^7] Dataset available at https://huggingface.co/datasets/copenlu/llm-pct-tropes We use the following LLMs: Mistral, Mixtral, Zephyr, Llama 2, Llama 3, OLMo (see Appendix A for further details).
@@ 4.1 Variability Through Persona Assignment
To assess variability of biases found in the models when prompted under different settings (RQ1), we first look at the coarse-grained impact of assigning personas to the model, i.e., when certain demographic categories in Table 1 are added to the prompt, in the closed setting. The overall results on the PCT across all models are provided in Fig. 2. As is visible in the plot, the responses of the models can change substantially when prompted with different personas, resulting in a change in their position on the PCT plot. One can also observe the variance of output of the different models under these conditions: Llama 3 and Zephyr’s answer positions change substantially based on the persona assigned to them in the prompt, especially when the category is political orientation. For example, in Fig. 3 we see that Mixtral can be pushed towards generating far right or far left stances simply by supplying the respective demographic in the prompt. Other models, such as OLMo and Llama 2, are less affected by demographic prompts, pointing to their steerability (Liu et al., 2024). We show standard deviations across responses, quantifying the impact of this further, in Appendix C, Fig. 16.
@@
Figure 5: Robustness comparison between open vs closed for Llama 3 for both the base case (i.e. no demographic prompting) and when the political orientation is “far right”.
@@
In order to quantify whether or not demographic features have a significant impact on placement on the PCT, we perform ordinary least-squares (OLS) regression on the outcomes of the PCT. OLS regression fits a simple linear model to the data, where significance can be measured based on the coefficients estimated from the population samples (in this case, varying the instructions for each individual demographic value). For this, we use the outcome $x$ and $y$ coordinates as dependent variables, and the demographic features encoded as a categorical variable for the independent variable. The results of this for the closed-form setting are given in Fig. 4.
@@ 4.2 Robustness Analysis
Next, we analyze the robustness of LLM stances in the presence of demographic features by looking at differences in closed-form stances and open ended stances, extending the work of Röttger et al. (2024) (RQ2). Fig. 5 shows responses from Llama 3 across the 62 PCT propositions for the open-ended and closed-form settings with no demographic based prompting (left side, which we denote as base case) as well as when prompting with the demographic “far right” (right side, comparison for other models can be seen in Appendix C). We see that Llama 3 tends to have high agreement with the propositions in the closed setting while disagreing more in the open setting. We also observe that in the open setting the model more often either a) refuses to answer, or b) tends to output a neutral stance. This is in line with Röttger et al. (2024) who show that models tend to shift their choices in the open setting. However, the difference between open-ended and closed-form responses is even more pronounced when introducing the “far right” demographic into the prompt. We find that there is stronger disagreement in this setting, where the responses disagree 90% of the time.
Given this, we demonstrate how these variations are systematic across models and settings by showing the average total variation distance (TVD) between the open-ended and closed-form responses of each model across demographic categories in Fig. 6. TVD measures the sum of absolute differences (i.e. L1 distance) between the probability mass for each set of responses. In other words, each response for each prompt is transformed into a vector of probabilities $\mathbf{p}_{x}=[P(\text{Agree}),P(\text{Disagree}),P(\text{None})]$, and the average TVD is calculated as
@@
We observe from Fig. 6 that the TVD tends to be higher for both right leaning demographics, demonstrating that models show high variation between the closed-form and open ended prompts when instructed to produce right-leaning political views. In contrast, left leaning demographics tend to have fewer differences across all models with the exception of OLMo, showing that prompting for far left positions is robust across prompting settings and potentially reflecting the observed left-leaning default stance among LLMs demonstrated in previous work Röttger et al. (2024); Hartmann et al. (2023). Finally, we see surprising results with OLMo, which expresses moderate TVD across all demographic categories, demonstrating that demographic features in the prompt have little impact on variations between the open domain and closed-form instruction formats. This echoes the results seen in Fig. 2, where demographic based prompts have little effect on OLMo.

### v2

# LLM Tropes: Revealing Fine-Grained Values and Opinions in Large Language Models

@@ Abstract
Uncovering latent values and opinions embedded in large language models (LLMs) can help identify biases and mitigate potential harm. Recently, this has been approached by prompting LLMs with survey questions and quantifying the stances in the outputs towards morally and politically charged statements. However, the stances generated by LLMs can vary greatly depending on how they are prompted, and there are many ways to argue for or against a given position. In this work, we propose to address this by analysing a large and robust dataset of 156k LLM responses to the 62 propositions of the Political Compass Test (PCT) generated by 6 LLMs using 420 prompt variations. We perform coarse-grained analysis of their generated stances and fine-grained analysis of the plain text justifications for those stances. For fine-grained analysis, we propose to identify tropes in the responses: semantically similar phrases that are recurrent and consistent across different prompts, revealing natural patterns in the text that a given LLM is prone to produce. We find that demographic features added to prompts significantly affect outcomes on the PCT, reflecting bias, as well as disparities between the results of tests when eliciting closed-form vs. open domain responses. Additionally, patterns in the plain text rationales via tropes show that similar justifications are repeatedly generated across models and prompts even with disparate stances.
LLM Tropes: Revealing Fine-Grained Values and Opinions in
Large Language Models
Dustin Wrightthanks:   denotes equal contribution      Arnav Arora11footnotemark: 1      Nadav Borenstein      Srishti Yadav Serge Belongie      Isabelle Augenstein University of Copenhagen {dw, aar, nb, srya, s.belongie, augenstein}@di.ku.dk
@@ 1 Introduction
First, recent work has shown that the responses of LLMs to survey questions depend highly on the phrasing of the question and the format of the answer (Wang et al., 2024; Röttger et al., 2024; Motoki et al., 2024), calling for a more robust evaluation setup for surfacing values embedded in language models. Second, when provided with different personas based on demographic characteristics, LLMs can reflect the social and political biases of the respective demographics Argyle et al. (2023), highlighting the need for disentangling the opinions embedded into LLMs and their variation when prompted with demographics. Such efforts also aid in aligning language models for different populations and cultures and prevent jailbreaking of LLMs Wei et al. (2023). Lastly, these evaluations focus primarily on quantifying stances towards the survey questions, ignoring justifications and explanations for those decisions. Revealing patterns in such data could offer a naturalistic way to express latent values and opinions in LLMs.
To address these shortcomings, we conduct a large-scale study eliciting 156,240 responses to the 62 propositions of the Political Compass Test (PCT) across 6 LLMs and 420 prompt variations. These prompt variations span different demographic personas: Age, Gender, Nationality, Political Orientation, and Class, as well as instruction prompts, in order to provide a robust set of data for analysis while disentangling demographic features. We propose to perform both coarse-grained analysis at the level of stances, allowing us to quantify political bias based on the PCT, as well as fine-grained analysis of the plain text open-ended justifications and explanations for those stances, allowing us to reveal latent values and opinions in the generated text. For fine-grained analysis, we propose to identify tropes in the responses: semantically similar phrases that are recurrent and consistent across different prompts, revealing patterns in the justifications that LLMs are prone to produce in different settings. An example from our dataset is given in Figure 1, where Llama 3 AI@Meta (2024) when prompted with two different demographic personas (being far left vs. far right) demonstrates both a tendency towards particular stances as well as particular lines of reasoning for those stances.
@@
  • We propose a naturalistic method for analysing bias in generated text through tropes, revealing the arguments which LLMs are likely to generate in different settings;
@@ 2 Related Work
Surfacing political biases embedded in NLP tools has been approached previously in the context of word embeddings Gordon et al. (2020) and masked language modeling (Schramowski et al., 2022; Stańczak et al., 2023). As these biases originate in the data the models are trained on Borenstein et al. (2024), there is also research on directly tying them back to training data (Feng et al., 2023). Recently, with more performant and coherent LLMs, extracting inherent biases have garnered more attention. This has been explored by eliciting responses from psychological surveys Miotto et al. (2022), opinion surveys Arora et al. (2023); Durmus et al. (2023); Santurkar et al. (2023), and personality tests (Rutinowski et al., 2023), finding that certain LLMs have a left-libertarian bias (Hartmann et al., 2023). There has also been research on examining the framing of the generations produced by LLMs (Bang et al., 2024), and on measuring biases across different social bias categories (Manerba et al., 2023). Recently, persona based evaluation of political biases has also been attempted to see if models can simulate the responses of populations (Liu et al., 2024; Hu and Collier, 2024; Jiang et al., 2022), including in languages other than English Thapa et al. (2023). However, impact of persona based evaluation remains underexplored.
Furthermore, there is research demonstrating the brittleness, values, and opinions of LLMs (Ceron et al., 2024). Our work is most similar to Röttger et al. (2024), who try to answer the question of how to meaningfully evaluate values and opinions in LLMs by way of the PCT. They demonstrate that eliciting open-ended responses from LLMs can lead to vastly different results than closed-form categorical selections, questioning the validity of biases found through such methods. They argue for more robust and domain specific evaluations. In order to further assess the sensitivity of these evaluations with respect to multiple personas, we conduct a systematic evaluation of variance across 156k generations. Further, considering the infeasability of analysing open-ended responses, we present a novel method for conducting such analysis through extraction of tropes from them.
@@ 3.1 Robust Dataset Generation
We generate a large dataset of responses to the PCT by varying two aspects of the prompt for each proposition: one aspect controlling the demographic category, and one aspect controlling the generation instruction. We improve over previous work, which has looked at robustness by varying the instructions or the order of appearance of options in the prompt Röttger et al. (2024); Hartmann et al. (2023); Motoki et al. (2024), by also varying across demographic features and analyzing the plain text justifications for the provided stances. As such, we adopt a general prompt template of the following form which we subsequently fill for the response to be generated: Prompt Template: Please follow all instructions from the perspective of a person who is {demographic}. Please write all responses in English. {instruction} where the demographic is selected from Table 1 and the instruction is selected from either Table 3 or Table 4 (see § A.1). We use 21 different demographic options, 20 instruction variations, and there are 62 PCT propositions, resulting in 26,040 responses per model that we study.
We cover both breadth and depth of demographic categories, covering 5 types of demographics and 3-5 values for each. For the instruction, prior work has shown that responses of an LLM change when constraints are put on the answer generation format (Röttger et al., 2024). Therefore, we use two settings: 1) open-ended generation, where no constraint is put on the model in terms of choosing a particular option; and 2) closed form generation, where the model is explicitly prompted to choose one of the listed options. In the closed form generation setting, the model is prompted to choose a stance towards the position based on the listed responses in the PCT survey: Strongly Agree, Agree, Disagree, Strongly Disagree. Additionally, we prompt the model to output an explanation for the selection. In the open-ended generation setting, the model is prompted with an open ended-prompt with no additional constraints. To further conduct coarse-grained analysis of the responses from the open-ended generation setting, including alignment between the open and closed settings, we categorise the open-ended responses of the LLMs into the selection options from the closed setting post-hoc using a Mistral-Instruct-v0.3 model (87% accuracy on a held out test set manually annotated by 3 annotators, see § A.7 for more details).
@@ 3.2 Tropes Extraction
Measuring categorical stances generated towards the PCT propositions provides coarse grained information about values and opinions by allowing one to quantify the political alignment of a given model/prompt combination. However, this coarse-grained information disregards the plain text justifications and explanations a model is likely to generate with respect to the propositions, which may reveal latent values and opinions not measured by the stances. In practice, users interact with LLMs in a plain-text, open-ended fashion, which this fine-grained information reflects. As such, we propose to complement the use of categorical stances with analysis of tropes present in the plain text responses. [^4] Tropes have been well studied in multiple domains, including literature Miller (1991) and television Gala et al. (2020). We adopt the following definition of a trope: A theme or motif which has the following properties: It is recurrent, i.e., appears frequently in the responses; and it is consistent, i.e., the statements which represent the trope can be grounded in a single abstract concept. Additionally, we focus specifically on tropes which convey a justification or explanation for the stances generated toward the political propositions. An example of such tropes is given in Figure 1. To find tropes, we propose to: 1) generate many responses for each proposition under different conditions (§ 3.1); 2) cluster individual sentences based on their semantic similarity, and (§ 3.2.1); 3) distill the semantic clusters to single high-level concepts, filtering those clusters which do not contain justifications or explanations relevant to the stance (§ 3.2.2).
@@ 3.2.2 Distilling Tropes
We remove outliers, clusters with fewer than 10 sentences or very large clusters that contain more than half of the sentences in $D^{\mathcal{P}}_{(\text{sup}/\text{opp})}$, filtering out 95% of the original $570k$ sentences in the dataset. We then distil each cluster to its main concept via the clusters’ centroids. For a given cluster $\mathcal{C}=\{\mathbf{e}^{1},...,\mathbf{e}^{|\mathcal{C}|}\}$, we first compute its Euclidean centre point $\mathbf{c}\in\mathbb{R}^{d}$, where $c_{i}=\frac{1}{|\mathcal{C}|}\sum_{\mathbf{e}\in\mathcal{C}}e_{i}$. Then, we find the cluster’s member $\hat{\mathbf{e}}$ that is the nearest to $\mathbf{c}$, that is, $\hat{\mathbf{e}}=\text{argmin}_{\mathbf{e}\in\mathcal{C}}\lVert\mathbf{e}-\mathbf{c}\lVert^{2}_{2}$. Mapping back the vectors $\hat{\mathbf{e}}$ to their sentences, we now have two sets of trope candidates for the proposition $P$: $T^{\mathcal{P}}_{\text{sup}}$ and $T^{\mathcal{P}}_{\text{opp}}$, one candidate for each cluster discovered by DBSCAN. However, not every candidate is a trope; many of the sentences in $T^{\mathcal{P}}_{(\text{sup}/\text{opp})}$ do not contain any argument or a justification relevant to the proposition. Such non-tropes are sentences such as “I agree with the proposition”, or “I believe that the potential benefits outweigh the challenges”. We use an LLM to filter out non-tropes from tropes, prompting it to classify if a trope candidate contains an explicit justification for agreeing or disagreeing with the proposition. See Appendix § A.3 for details.
@@ 4 Analysis
Our analysis centers around four research questions addressing the three shortcomings of previous work described in § 1: RQ1: How do demographic-based prompts impact LLM survey responses? RQ2: Do categorical surveys administered to LLMs elicit robust results over diverse prompts? RQ3: What tropes do LLMs produce in response to the PCT? RQ4: Do variations in categorical survey responses reflect variations in the tropes? To answer these questions, we generate a large dataset of responses to the PCT as described in § 3.1, eliciting 26,040 responses for 6 different language models (156,240 responses total). [^7] Dataset available at https://huggingface.co/datasets/copenlu/llm-pct-tropes We use the following LLMs: Mistral, Mixtral, Zephyr, Llama 2, Llama 3, OLMo (seeApp. A for further details).
@@ 4.1 Variability Through Persona Assignment
To assess variability of biases found in the models when prompted under different settings (RQ1), we first look at the coarse-grained impact of assigning personas to the model, i.e., when certain demographic categories in Table 1 are added to the prompt, in the closed setting. The overall results on the PCT across all models are provided in Figure 2. As is visible in the plot, the responses of the models can change substantially when prompted with different personas, resulting in a change in their position on the PCT plot. One can also observe the variance of output of the different models under these conditions: Llama 3 and Mixtral’s answer positions change substantially based on the persona assigned to them in the prompt, especially when the category is political orientation. For example, in Figure 3 we see that Mixtral can be pushed towards generating far right or far left stances simply by supplying the respective demographic in the prompt. Other models, such as OLMo and Llama 2, are less affected by demographic prompts, pointing to their steerability (Liu et al., 2024). We show standard deviations across responses, quantifying the impact of this further, in App. C, Figure 16.
@@
Figure 5: Robustness comparison between the open-domain prompts and closed-form prompts for Llama 3. On the left we show the base case with no demographic prompting and on the right we show the case for “far right”. Each bar represents one question on the PCT, and the colors indicate the distribution of responses to that question across instruction prompts (dark blue is strong agree, light blue is agree, light orange is disagree, dark orange is strong disagree, and grey is refusal to answer or taking a neutral stance).
@@
In order to quantify whether or not demographic features have a significant impact on placement on the PCT, we perform ordinary least-squares (OLS) regression on the outcomes of the PCT. For this, we use the outcome $x$ and $y$ coordinates as dependent variables, and the demographic features encoded as a categorical variable for the independent variable. The results of this for the closed-form setting are given in Figure 4.
@@ 4.2 Robustness Analysis
Previous work shows that the stances produced by closed-form instructions can greatly diverge from stances produced by open-ended generation (Röttger et al., 2024). Here, we explore how demographic features impact this disparity, and if certain demographics produce open-ended responses with stances which better reflect their forced-choice stances (RQ2). Figure 5 shows responses from Llama 3 across the 62 PCT propositions for the open-ended and closed-form settings with no demographic based prompting (left side, which we denote as base case) as well as when prompting with the demographic “far right” (right side, comparison for other models can be seen in App. C). We see that Llama 3 tends to have high agreement with the propositions in the closed setting while disagreing more in the open setting. We also observe that in the open setting the model more often either a) refuses to answer, or b) tends to output a neutral stance. This is in line with Röttger et al. (2024) who show that models tend to shift their choices in the open setting. However, the difference between open-ended and closed-form responses is even more pronounced when introducing the “far right” demographic into the prompt. We find that there is stronger disagreement in this setting, where the responses disagree 90% of the time.
Given this, we demonstrate how these variations are systematic across models and settings by showing the average total variation distance (TVD) between the open-ended and closed-form responses of each model across demographic categories in Figure 6. TVD measures the sum of absolute differences (i.e. L1 distance) between the probability mass for each set of responses. In other words, each response for each prompt is transformed into a vector of probabilities $\mathbf{p}_{x}=[P(\text{Agree}),P(\text{Disagree}),P(\text{None})]$, and the average TVD is calculated as
@@
We observe from Figure 6 that the TVD changes substantially with demographic. Most notably, left leaning demographics have much lower TVD across all models. This demonstrates that prompting for far left positions leads to less variation between different prompts, providing stronger evidence for a left-leaning default stance among LLMs as demonstrated in previous work (Röttger et al., 2024; Hartmann et al., 2023). Additionally, though OLMo is the least influenced by demographic features in the closed-form setting (see Figure 2), it has generally much higher TVD across all settings, demonstrating its higher sensitivity to change in output format compared to other models. Overall, from these results, we conclude that the variation in outcomes between prompt types is systematic with the exception of prompts using left-leaning demographics, which result in similar outcomes regardless of prompt type.